OpenStudy (anonymous):

Solve the equation. 8 ( r + 1 ) = - 2 ( r - 44 )

OpenStudy (anonymous):

r=8

OpenStudy (anonymous):

4(r+1)=-r+44 4r+4=-r+44 5r=40 r=8

OpenStudy (anonymous):

8(r+1)=-2(r-44) 8r+8=-2r+88 (then you add 2r to -2r which cancels it out and add 2r to 8r) 10r+8=88 (then you minus 8 to 88) 10r=80 (finally you divide) r=8

OpenStudy (anonymous):

the reason why you added -2r and 2r and not -2r-2r is because you want to do the inverse operation.Also simplifying each step makes it shorter.